• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sarty

    Log entry time 07:50:12 on June 4,2000

    Entry number 43068

    keyword=shift summary

    Productive shift:
    ...pointing ref measurements for thcm=+155
    ...harp scan
    ...sieve run for HRSE at 14.07 deg
    ...pointing ref measurements for thcm=-155
    ...production running for thcm=-155: TOTAL TRIGS = 6.92 M
    .
    Comments:
    ...lots of rf trips (cavity in north linac). Beam position a little unstable always (for about a minute) when returning from rf trips.
    .
    ****************************************************
    Details below:
    .
    Start shift with Pointing Ref measurements for just-completed production
    running (thcm=+155, phi=0):
    .....00:15-00:25: Run 1753: 10 cm dummy (5 uA, unrastered)
    .....00:30-00:35: Run 1755: 15 cm dummy (5 uA, unrastered) (Run 1754 junk)
    .....00:40-00:45: Run 1756: 4 cm dummy (5 uA, unrastered)
    .....00:50-00:55 Run 1758: carbon (5 uA, unrastered) (Run 1757 junk)
    .
    .....01:05-01:06: Run 1759: harp scan (Harp scan 924)
    .
    Changing HRSH angle to 49.72 degrees for Thcm=-155 deg (phi=180):
    .
    01:19-01:36: Run 1760: sieve slit run in HRSE, carbon (while HRSH moves angle)
    .
    Pointing Ref for Thcm=-155 deg:
    .
    01:50-01:55: Run 1761: carbon (20 uA, unrastered)
    02:00-02:05: Run 1762: 4 cm dummy (5 uA, unrastered)
    02:10-02:14: Run 1763: 15 cm dummy (5 uA, unrastered)
    02:17-02:20: Run 1764: 10 cm dummy (5 uA, unrastered)
    .
    Production running for thcm=-155 deg:
    .
    02:35: Run 1765: Production at 47.5 uA.......................................1.25 M trigs
    03:15: Run 1766: Production at 47.5 uA.......................................0.77 M trigs
    .
    03:45 - beam off: many rf trips; stopping to look for fix (beam was off for last about 10 minutes of Run 1766)
    04:30 - MCC bringing beam back to us which "might be stable"
    .
    04:33 - Run 1767: junk - beam didn't last; have to bypass a cavity in north linac.
    04:45 - Run 1768: trying again (47.8 uA)......................................1.25 M trigs
    05:38 - Run 1769: prod run (47 uA)..............................................1.25 M trigs
    06:25 - Run 1770: prod run (47 uA)...............................................1.25 M trigs
    07:05 - Run 1771: prod run (46.5 uA).............................................1.15 M trigs
    .
    07:47 - beam off for rf trips plus making QE measurement